Jefferson Lab HALL A Logbook for March 07 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

    
    

    Friday

  • 197674 03/30/07 17:53 A. Camsonne Additionnal delay on S1 2L and 3L
  • 197673 03/30/07 15:22 R. Michaels trigger prep, part II
  • 197672 03/30/07 14:50 R. Michaels hatsv3, hatsv10
  • 197671 03/30/07 13:54 atrig Trigger-Download
  • 197670 03/30/07 13:53 atrig Trigger-Download
  • 197669 03/30/07 13:52 atrig Trigger-Download
  • 197668 03/30/07 13:52 atrig Trigger-Download
  • 197667 03/30/07 13:50 atrig Trigger-Download
  • 197666 03/30/07 13:48 atrig Trigger-Download
  • 197665 03/30/07 13:47 atrig Trigger-Download
  • 197664 03/30/07 13:41 atrig Trigger-Download
  • 197661 03/30/07 13:39 atrig Trigger-Download
  • 197660 03/30/07 13:38 atrig Trigger-Download
  • 197659 03/30/07 13:37 atrig Trigger-Download
  • 197656 03/30/07 13:33 atrig Trigger-Download
  • 197652 03/30/07 13:15 atrig Trigger-Download
  • 197651 03/30/07 13:13 atrig Trigger-Download
  • 197650 03/30/07 08:21 fansler Q2 Cryo valves
  • Thursday

  • 197639 03/29/07 16:56 Makis P Right S1 HVs reset to E04-018 values
  • 197616 03/29/07 14:32 fansler N2 liquid level meters
  • 197609 03/29/07 13:15 A. Camsonne Right arm S1 OR and S2m OR in scaler
  • 197604 03/29/07 12:53 A. Camsonne Left S1 OR and S2 OR in scaler
  • 197591 03/29/07 10:14 fansler N2 to hall A
  • Wednesday

  • 197590 03/28/07 17:04 R. Michaels trigger adjustments part I
  • 197586 03/28/07 16:32 atrig Trigger-Download
  • 197585 03/28/07 16:30 fansler Liquid Level Meters
  • 197584 03/28/07 16:25 atrig Trigger-Download
  • 197583 03/28/07 16:24 atrig Trigger-Download
  • 197582 03/28/07 16:23 atrig Trigger-Download
  • 197581 03/28/07 16:21 atrig Trigger-Download
  • 197580 03/28/07 16:20 atrig Trigger-Download
  • 197579 03/28/07 16:20 atrig Trigger-Download
  • 197577 03/28/07 16:18 atrig Trigger-Download
  • 197576 03/28/07 16:17 atrig Trigger-Download
  • 197575 03/28/07 16:08 atrig Trigger-Download
  • 197574 03/28/07 16:03 atrig Trigger-Download
  • 197573 03/28/07 16:01 atrig Trigger-Download
  • 197572 03/28/07 15:47 atrig Trigger-Download
  • 197571 03/28/07 15:45 atrig Trigger-Download
  • 197570 03/28/07 15:42 atrig Trigger-Download
  • 197569 03/28/07 15:41 atrig Trigger-Download
  • 197568 03/28/07 15:41 atrig Trigger-Download
  • 197567 03/28/07 15:39 atrig Trigger-Download
  • 197566 03/28/07 15:38 atrig Trigger-Download
  • 197565 03/28/07 15:37 atrig Trigger-Download
  • 197560 03/28/07 14:49 atrig Trigger-Download
  • 197559 03/28/07 14:49 atrig Trigger-Download
  • 197558 03/28/07 14:45 atrig Trigger-Download
  • 197557 03/28/07 14:44 atrig Trigger-Download
  • 197555 03/28/07 14:38 atrig Trigger-Download
  • 197554 03/28/07 13:23 A. Camsonne Backup script directory in save-28Mar07/
  • 197552 03/28/07 11:24 atrig Trigger-Download
  • Tuesday

  • 197551 03/27/07 15:11 M. Stevens Spectrometer magnets functionally test o.k.
  • 197550 03/27/07 13:30 spiegel Hall A post -beam Systems check out
  • 197549 03/27/07 13:13 fansler Cryo Valves
  • 197548 03/27/07 09:01 folts thec / Hall A to restricted access
  • 197547 03/27/07 06:41 folts tech/High rad levels
  • Monday

  • 197546 03/26/07 15:08 M Jones E06-007 , Lead cross section experiment has ended
  • 197545 03/26/07 14:17 Holmstrom Target warm Up and HRS off
  • 197542 03/26/07 11:08 A. Camsonne Hall A radiation level and raster mystery
  • 197544 03/26/07 10:45 reinhold timestamps from webinterface entries appear 1 hour early
  • 197540 03/26/07 10:42 A. Camsonne Pedestal subtraction off for pedestal check
  • 197543 03/26/07 10:40 Reinhold Harp scan history: beam spot was degrading
  • 197539 03/26/07 09:27 folts High Rads in the Hall
  • 197538 03/26/07 08:20 B. Hahn Owl Shift summary
  • 197519 03/26/07 04:57 B Hahn run 2005
  • 197518 03/26/07 04:56 M Jones Kin07 Bi General tools screen
  • 197514 03/26/07 04:34 J. Lopez Hall Access to reboot Right NMR
  • 197513 03/26/07 03:31 J. Lopez No conecction to hareboot27
  • 197512 03/26/07 02:56 J. Lopez D1 Lead A Helium Flow is flat
  • 197511 03/26/07 02:50 J. Lopez Result for harp scan after adjustments
  • 197510 03/26/07 02:26 J. Lopez Results from Harp Scan
  • 197509 03/26/07 02:05 J Lopez Moved to empty target for a harp scan
  • 197507 03/26/07 01:51 J. Lopez SLI screen
  • 197506 03/26/07 01:37 J. Lopez Strip Charts
  • 197503 03/26/07 01:31 J. Lopez Target Ladder Camera is not working
  • 197502 03/26/07 01:13 Adaq Raster at Bi (Kin 5) Run 2002
  • 197499 03/26/07 01:04 J. Lopez Hall A General Tools
  • Sunday

  • 197491 03/25/07 23:57 bogdanw shift summary
  • 197483 03/25/07 23:06 bogdanw HV screens
  • 197490 03/25/07 22:56 reinhold Re: SLI is larger; and you're hitting also much more now
  • 197474 03/25/07 21:23 bogdanw SLI is larger
  • 197459 03/25/07 19:21 bogdanw energy spread
  • 197458 03/25/07 19:11 bogdanw stripcharts
  • 197455 03/25/07 19:02 Ramesh Bi target
  • 197451 03/25/07 18:42 Ramesh Target to empty
  • 197450 03/25/07 18:37 bogdanw setting for kin#3
  • 197438 03/25/07 16:01 X. Zheng day shift summary
  • 197436 03/25/07 15:49 S.Abrahamyan Prescale change for Kin3.
  • 197435 03/25/07 15:43 X. Zheng Tools screen
  • 197432 03/25/07 15:41 X. Zheng strip charts
  • 197424 03/25/07 14:03 X. Zheng stripchart - energy spread now slightly larger
  • 197423 03/25/07 13:56 X. Zheng SLI screen
  • 197420 03/25/07 13:48 Shahinyan Target
  • 197417 03/25/07 13:22 Reinhold spot: looks like a broken target: already seen earlier
  • 197416 03/25/07 13:11 Reinhold spot for 1954, first Bi run
  • 197413 03/25/07 13:02 Reinhold spot: looks like a broken target
  • 197411 03/25/07 12:39 X. Zheng script chart - glitch in left dipole field
  • 197408 03/25/07 12:11 X. Zheng left dipole status
  • 197402 03/25/07 11:15 X. Zheng BCM screen and stripchart
  • 197401 03/25/07 11:04 X. Zheng magnet stripchart
  • 197400 03/25/07 11:02 X. Zheng Left dipole seems okay now (keep monitoring)
  • 197398 03/25/07 10:42 X. Zheng Left Dipole power supply
  • 197397 03/25/07 10:41 X. Zheng more on left dipole
  • 197396 03/25/07 10:36 X. Zheng Left magnet alarm
  • 197394 03/25/07 10:04 X. Zheng SLI screen
  • 197391 03/25/07 09:53 M Jones Hall A General tools
  • 197390 03/25/07 09:41 X. Zheng stripchart -- jump in HacL_D1_P0rb
  • 197385 03/25/07 08:51 X. Zheng spot++
  • 197383 03/25/07 08:31 S.Abrahamyan Raster size for run 1961
  • 197382 03/25/07 08:22 Shahinyan Target move
  • 197380 03/25/07 08:08 Shahinyan Target Status
  • 197379 03/25/07 08:04 Jon Miller Shift Summary
  • 197378 03/25/07 08:03 Jon Miller Strip charts
  • 197377 03/25/07 08:02 Jon Miller General Tools
  • 197352 03/25/07 04:13 Jon Miller Strip Charts
  • 197351 03/25/07 04:12 Jon Miller General Tools
  • 197353 03/25/07 03:16 reinhold possible raster explanation
  • 197342 03/25/07 01:07 J Lopez Lead (Emiss spectrum) KIN 12
  • 197335 03/25/07 00:36 Jon Miller Strip Chart Fix
  • 197332 03/25/07 00:24 reinhold possible raster explanation
  • 197331 03/25/07 00:13 Jon Miller Strip Charts
  • 197330 03/25/07 00:13 Jon Miller General Tools
  • 197329 03/25/07 00:06 Shahinyan Shift Summar
  • 197328 03/25/07 00:04 reinhold possible raster explanation
  • Saturday

  • 197324 03/24/07 23:57 reinhold possible raster explanation
  • 197323 03/24/07 23:54 reinhold possible raster explanation
  • 197322 03/24/07 23:52 reinhold possible raster explanation
  • 197321 03/24/07 23:47 Reinhold possible raster explanation
  • 197318 03/24/07 22:26 J Lopez Lead target in kinematics 4
  • 197307 03/24/07 20:48 Reinhold spot raster at 40 uA (back to bad)
  • 197302 03/24/07 19:57 reinhold spot raster for run 1932 (looks much better)
  • 197298 03/24/07 19:43 reinhold moved to lead target
  • 197295 03/24/07 19:30 Ramesh Ready for Kin 04
  • 197294 03/24/07 19:07 Reinhold investigate raster history
  • 197292 03/24/07 18:56 Reinhold move target to carbon for harp scan
  • 197290 03/24/07 18:46 Adaq Start kinematic #11 run 1916
  • 197289 03/24/07 18:42 A. Camsonne T1 and T3 divided by charge for Kin11 and Kin 6
  • 197293 03/24/07 18:00 J Lopez Lead is still alive
  • 197280 03/24/07 16:21 reinhold spot (2nd half of figure)
  • 197279 03/24/07 16:16 Reinhold spot raster run 1926
  • 197278 03/24/07 16:03 Reinhold Target status at beginning of shift: Pb target
  • 197276 03/24/07 15:59 X. Zheng day shift summary
  • 197275 03/24/07 15:52 X. Zheng SLI screen
  • 197274 03/24/07 15:50 S.Abrahamyan Raster size for run 1926
  • 197273 03/24/07 15:48 X. Zheng helium flow chart
  • 197270 03/24/07 15:45 X. Zheng Current monitor and SLI stripchart
  • 197263 03/24/07 13:23 S.Abrahamyan raster size run 1922
  • 197262 03/24/07 13:18 X.Zheng current monitor
  • 197261 03/24/07 13:17 S.Abrahamyan raster size
  • 197260 03/24/07 13:16 R.Lindgren Strip Charts
  • 197257 03/24/07 12:26 R. Lindgren Beam energy spread strip tool
  • 197254 03/24/07 12:20 R. Lindgren Cryo Levels
  • 197253 03/24/07 12:17 R.Lindgren Hall A Tools
  • 197249 03/24/07 11:31 X. Zheng no beam for 1/2 hour
  • 197245 03/24/07 10:41 Aniol rate increases 1872/1871
  • 197244 03/24/07 10:18 S.Abrahamyan Target Solid3
  • 197240 03/24/07 08:46 S.Abrahamyan Target empty
  • 197235 03/24/07 08:02 R. Michaels owl summary
  • 197231 03/24/07 07:22 R. Michaels ps1-4 adjustment
  • 197217 03/24/07 04:51 R. Michaels TS0 hangup, restart CODA
  • 197215 03/24/07 04:25 R. Michaels MCC, halfwave plate
  • 197214 03/24/07 04:09 B Hahn run type
  • 197207 03/24/07 03:18 R. Michaels rastXY dependence on trigger
  • 197194 03/24/07 01:03 R. Michaels beam current dependent bgr ?
  • 197191 03/24/07 00:38 R. Michaels kin 11, general tools
  • 197188 03/24/07 00:20 Reinhold shift summary
  • 197186 03/24/07 00:14 Reinhold try smaller raster on C
  • 197182 03/24/07 00:01 reinhold talked to RC
  • Friday

  • 197179 03/23/07 23:58 reinhold try carbon with this raster
  • 197178 03/23/07 23:57 reinhold move to carbon target
  • 197175 03/23/07 23:34 reinhold halog messes with the figures
  • 197174 03/23/07 23:32 reinhold spot again (is the target still OK?)
  • 197187 03/23/07 23:20 M Jones T3 rates ( Time submitted 12:02)
  • 197171 03/23/07 23:13 reinhold I'm confused about the raster
  • 197168 03/23/07 22:53 reinhold spot comparison 1886 1888
  • 197165 03/23/07 22:28 reinhold changed raster to 6.0 x 6.4
  • 197155 03/23/07 21:47 reinhold run 1884 is pulser run
  • 197152 03/23/07 21:39 reinhold changed raster to 7.3 x 5.2
  • 197150 03/23/07 21:29 reinhold current raster screen
  • 197149 03/23/07 21:28 reinhold spot comparison 1882 vs 1864
  • 197147 03/23/07 21:10 reinhold spectrometer motion
  • 197146 03/23/07 21:08 reinhold Konrad calls: sees T3 increase
  • 197145 03/23/07 20:55 reinhold spot
  • 197140 03/23/07 20:12 reinhold Raster at target run 1877
  • 197138 03/23/07 20:08 reinhold beam current & SLI strip tools
  • 197125 03/23/07 17:55 reinhold beam current and SLI strip tools
  • 197122 03/23/07 17:31 reinhold sli screen 6e-5
  • 197120 03/23/07 17:29 reinhold ask for 40 uA
  • 197117 03/23/07 17:17 reinhold ask for 20 uA
  • 197114 03/23/07 17:07 reinhold moved to lead Pb target
  • 197113 03/23/07 17:07 reinhold harp scan
  • 197111 03/23/07 17:03 reinhold spot at 7.3 x 5.8
  • 197110 03/23/07 17:01 reinhold spot at 7.3 x 5.8
  • 197109 03/23/07 16:57 reinhold spot with 6.5 x 6.4 raster
  • 197105 03/23/07 16:31 reinhold raster was a bit to large
  • 197104 03/23/07 16:28 reinhold change raster to 6.5 by 6.4
  • 197103 03/23/07 16:25 reinhold moved carbon target in
  • 197101 03/23/07 16:18 reinhold current at 40 uA now
  • 197099 03/23/07 16:10 reinhold Ask MCC for 20 uA
  • 197097 03/23/07 15:02 LeRose Shift Summary (16:02)
  • 197095 03/23/07 14:33 LeRose Beam
  • 197093 03/23/07 14:10 LeRose Beam Back (sort of)
  • 197092 03/23/07 14:09 Ole Hansen Move target to lead
  • 197091 03/23/07 13:12 Ole Hansen Move target to empty
  • 197090 03/23/07 12:36 LeRose MCC running Crest
  • 197089 03/23/07 12:17 LeRose IOC strip chart
  • 197088 03/23/07 12:16 LeRose Tools/Beam
  • 197087 03/23/07 12:05 R. Michaels rastxy restored
  • 197086 03/23/07 11:53 LeRose Beam Almost Back
  • 197085 03/23/07 09:15 LeRose No Beam
  • 197083 03/23/07 09:11 LeRose Spot++
  • 197076 03/23/07 08:04 Cornejo, J.C Owl shift summary
  • 197072 03/23/07 07:21 Cornejo, J.C Beam spread screenshot
  • 197071 03/23/07 07:21 Cornejo, J.C Strip graphs
  • 197068 03/23/07 07:20 Cornejo, J.C Current monitor
  • 197067 03/23/07 07:19 Cornejo, J.C Hall A Tools
  • 197066 03/23/07 07:11 Cornejo, J.C raster spectra run 1859
  • 197065 03/23/07 07:01 Cornejo, J.C spot++ run 1860
  • 197062 03/23/07 06:50 Cornejo, J.C Beam energy spread
  • 197061 03/23/07 06:49 Cornejo,J.C. SLI screenshot
  • 197060 03/23/07 06:45 Cornejo, J.C Beam current screenshot
  • 197059 03/23/07 06:37 Cornejo, J.C Back at 40uA beam for kinematics 09
  • 197056 03/23/07 06:32 Cornejo, J.C raster spectra run 1858
  • 197055 03/23/07 06:29 Cornejo, J.C Raster spectra run 1857
  • 197054 03/23/07 06:27 Cornejo, J.C Runs 1857 and 1858 are kinematics 9
  • 197051 03/23/07 06:16 Cornejo, J.C Spot++ after spectrometer change.q
  • 197049 03/23/07 06:12 Cornejo, J.C Receiving 10uA ( warm up )
  • 197048 03/23/07 06:07 Cornejo, J.C Beam Spread screenshots
  • 197047 03/23/07 06:06 Cornejo, J.C Strips screenshot
  • 197046 03/23/07 06:05 Cornejo, J.C Hall A Tools
  • 197045 03/23/07 06:05 Cornejo, J.C Right arm moved to 30.902
  • 197044 03/23/07 06:04 Chiranjib target moved to Pb
  • 197043 03/23/07 06:04 Cornejo, J.C Raster spectra 1856
  • 197042 03/23/07 05:32 Cornejo, J.C Changing Right spectrometer angle
  • 197041 03/23/07 05:32 Cornejo, J.C Harp scan results
  • 197040 03/23/07 05:11 Chiranjib target moved to empty
  • 197038 03/23/07 04:57 Cornejo, J.C Ramping current down slowly
  • 197031 03/23/07 04:12 Cornejo, J.C SLI screenshot
  • 197028 03/23/07 03:35 Cornejo, J.C Spot++ run 1852
  • 197023 03/23/07 03:05 Cornejo, J.C raster spectra run 1850
  • 197022 03/23/07 02:59 Cornejo, J.C Beam Spread screenshot
  • 197021 03/23/07 02:58 Cornejo, J.C Strip tool screenshot
  • 197020 03/23/07 02:58 Cornejo, J.C Current screenshot
  • 197019 03/23/07 02:57 Cornejo, J.C Hall A Tools screenshot
  • 197014 03/23/07 02:25 Cornejo, J.C Raster spectra run 1848
  • 197011 03/23/07 01:55 Cornejo, J.C Raster spectra run 1846
  • 197008 03/23/07 01:37 Cornejo, J.C Spot++ run 1846
  • 197003 03/23/07 00:50 Cornejo, J.C Energy spread screenshot
  • 197002 03/23/07 00:48 Cornejo, J.C Strip snapshot
  • 197001 03/23/07 00:47 Cornejo, J.C Current snapshot
  • 197000 03/23/07 00:47 Cornejo, J.C Hall A Tools
  • 196995 03/23/07 00:08 Eric Voutier Shift summary
  • Thursday

  • 196989 03/22/07 23:41 Adaq Emiss improved energy resolution
  • 196988 03/22/07 23:37 Ramesh Hall A tools
  • 196987 03/22/07 23:34 Ramesh Target screen
  • 196977 03/22/07 21:35 R. Michaels access, chk cable, relink spot++
  • 196976 03/22/07 21:10 R. Michaels the story of spot++
  • 196970 03/22/07 19:40 Ramesh spot chek again
  • 196969 03/22/07 19:31 Ramesh spot check on C, may be scale problem
  • 196964 03/22/07 18:55 Ramesh Spot unusual
  • 196960 03/22/07 15:58 Higinbotham Shift Summary
  • 196955 03/22/07 14:36 A. Camsonne Right arm IOC reboot
  • 196958 03/22/07 14:23 A. Camsonne EPICS start of run test
  • 196956 03/22/07 13:36 A. Camsonne Run 1829 test run after right arm reboot
  • 196952 03/22/07 08:53 Ole Hansen Target to empty
  • 196950 03/22/07 08:13 Reinhold Shift Summary
  • 196944 03/22/07 06:41 reinhold MCC can't do more than 30 uA
  • 196940 03/22/07 06:27 reinhold ask MCC for 30 uA
  • 196937 03/22/07 06:18 reinhold MCC calls: might not need access
  • 196935 03/22/07 06:11 reinhold MCC call
  • 196932 03/22/07 05:49 Chiranjib raster RUN #1821
  • 196931 03/22/07 05:47 reinhold beam current and sli strip tool
  • 196930 03/22/07 05:44 reinhold performed shift checklist
  • 196927 03/22/07 05:35 Reinhold General Tools screen
  • 196926 03/22/07 05:32 reinhold spot check run 1821
  • 196925 03/22/07 05:20 Chiranjib raster RUN #1820
  • 196920 03/22/07 04:28 Chiranjib raster at target RUN # 1819
  • 196919 03/22/07 04:10 reinhold SLI screen: dE~4e-5
  • 196918 03/22/07 04:09 reinhold beam current and sli screens
  • 196915 03/22/07 03:51 Chiranjib raster at target RUN # 1818
  • 196912 03/22/07 03:05 reinhold beam current and SLI last hour
  • 196911 03/22/07 03:03 Chiranjib raster at target RUN #1817
  • 196910 03/22/07 03:00 Chiranjib raster at target RUN #1816
  • 196909 03/22/07 02:56 CHiranjib raster at target RUN # 1815
  • 196908 03/22/07 02:47 reinhold spot check run 1816
  • 196903 03/22/07 01:56 reinhold energy spread at 40 uA ~3.5e-5
  • 196902 03/22/07 01:46 reinhold spot at 40 uA
  • 196901 03/22/07 01:44 reinhold current monitor screen: target warmup history
  • 196899 03/22/07 01:39 reinhold ask for 40 uA
  • 196897 03/22/07 01:29 reinhold spot at 30 uA
  • 196896 03/22/07 01:28 reinhold spot at 20 uA
  • 196894 03/22/07 01:25 reinhold increase beam to 30 uA
  • 196892 03/22/07 01:17 Cornejo, J.C raster spectra run 1812
  • 196890 03/22/07 01:16 reinhold energh spread at 20 uA
  • 196889 03/22/07 01:14 reinhold ask for 20 uA
  • 196887 03/22/07 01:11 reinhold raster screen
  • 196886 03/22/07 01:11 reinhold energy spread at 10uA
  • 196885 03/22/07 01:07 reinhold spot with lead (Pb) at 10 uA
  • 196883 03/22/07 01:01 reinhold PB target in; ask for 10 uA
  • 196882 03/22/07 01:01 Chiranjib target moved to Pb
  • 196881 03/22/07 00:55 reinhold ask beam off for target motion
  • 196880 03/22/07 00:51 reinhold spot with carbon
  • 196878 03/22/07 00:45 reinhold MCC has problems going to 40 uA
  • 196876 03/22/07 00:43 Reinhold position feedback was off
  • 196875 03/22/07 00:40 Reinhold spot check
  • 196872 03/22/07 00:25 Chiranjib Target moved to C
  • 196871 03/22/07 00:21 Reinhold MCC ready with beam
  • 196870 03/22/07 00:18 Reinhold MCC calls: will send beam
  • 196868 03/22/07 00:05 Makis P Shift Summary
  • Wednesday

  • 196867 03/21/07 23:45 Mina Kat target moved to empty
  • 196866 03/21/07 22:55 Makis P No beam for 1 hr/Asked for acess
  • 196865 03/21/07 22:39 Makis P beam off/power loss acc racks
  • 196863 03/21/07 22:31 Makis P Peanuts Level
  • 196858 03/21/07 21:45 Makis P Current Monitor
  • 196857 03/21/07 21:44 Makis P Hall A Tools
  • 196856 03/21/07 21:19 Mina Kat IOC monitoring
  • 196849 03/21/07 20:02 Mina Kat spot++ run 1804
  • 196846 03/21/07 19:46 Mina Kat Target Charts
  • 196843 03/21/07 19:00 Mina Kat Energy spread monitoring
  • 196838 03/21/07 18:35 Makis P beam spread grows
  • 196833 03/21/07 18:01 Mina Kat spot++ run 1798 Pb+C#3
  • 196826 03/21/07 17:29 Mina Kat Beam raster vs Current run 1794
  • 196821 03/21/07 17:11 Mina Kat spot++ for run 1793 Pb+C#3
  • 196820 03/21/07 17:01 Makis P Raster parameter change (same area)
  • 196818 03/21/07 17:00 Mina Kat Target changed to Pb+C#3
  • 196810 03/21/07 16:28 Mina Kat target changed to C
  • 196809 03/21/07 16:26 Ramesh Raster check on C
  • 196803 03/21/07 16:07 Higinbotham Shift Summary
  • 196802 03/21/07 16:06 Higinbotham Raster Aspect Flipped
  • 196800 03/21/07 15:59 R. Michaels raster might be too wide
  • 196793 03/21/07 15:23 R. Michaels moving to Pb#3
  • 196782 03/21/07 08:52 R. Michaels spot++ shows a big peak
  • 196780 03/21/07 08:01 C. Fernandez Owl Shift Summary
  • 196774 03/21/07 06:53 C. Fernandez energy spread snapshot
  • 196773 03/21/07 06:52 C. Fernandez strip snapshot
  • 196772 03/21/07 06:52 C. Fernandez current snapshot
  • 196771 03/21/07 06:52 C. Fernandez Hall A general tools
  • 196763 03/21/07 02:39 Reinhold Target: moved from empty to carbon
  • 196762 03/21/07 02:38 C. Fernandez Beam on
  • 196761 03/21/07 01:24 C. Fernandez No beam
  • 196760 03/21/07 00:09 Reinhold Target status (cont.)
  • 196759 03/21/07 00:08 Reinhold Target status
  • Tuesday

  • 196757 03/20/07 23:59 E.Chudakov Swing shift summary
  • 196756 03/20/07 23:18 E.Chudakov Tools, beam energy spread
  • 196755 03/20/07 23:17 E.Chudakov MCC progress
  • 196754 03/20/07 22:52 E.Chudakov Harp scan
  • 196752 03/20/07 15:59 Higinbotham Shift Summary
  • 196751 03/20/07 15:38 Higinbotham Right HRS to 83.23 degrees
  • 196749 03/20/07 08:34 R. Michaels general tools
  • 196748 03/20/07 08:22 R. Michaels replay problems
  • 196745 03/20/07 08:10 Vignote raster plots, run 1771
  • 196744 03/20/07 08:03 C. Fernandez Owl Shift Summary
  • 196740 03/20/07 07:35 Cornejo, J.C Raster spectra run 1769
  • 196735 03/20/07 06:43 C. Fernandez energy spread snapshot
  • 196732 03/20/07 06:23 Cornejo, J.C Raster spectra run 1767
  • 196731 03/20/07 06:07 C. Fernandez energy spread snapshot
  • 196729 03/20/07 06:06 C. Fernandez strip snapshot
  • 196727 03/20/07 06:05 C. Fernandez current snapshot
  • 196726 03/20/07 06:05 C. Fernandez Hall A general tools snapshot
  • 196725 03/20/07 06:04 C. Fernandez spot++ snapshot
  • 196722 03/20/07 05:15 Cornejo, J.C raster spectra run 1765
  • 196717 03/20/07 04:22 Cornejo, J.C Raster spectra run 1763
  • 196716 03/20/07 04:18 C. Fernandez spot++ snapshot
  • 196713 03/20/07 04:01 Cornejo, J.C Raster spectra run 1761
  • 196710 03/20/07 03:16 Cornejo, J.C Raster spectra run 1759
  • 196709 03/20/07 03:15 C. Fenandez- Energy spread snapshot
  • 196708 03/20/07 03:15 C. Fernandez strip snapshot
  • 196707 03/20/07 03:14 C. Fernandez current snapshot
  • 196706 03/20/07 03:14 C. Fernandez Hall A general tool snapshot
  • 196703 03/20/07 02:59 C. Fernandez spot++ snapshot
  • 196698 03/20/07 01:50 C. Fernandez energy spread snapshot
  • 196697 03/20/07 01:49 C. Fernandez strip snapshot
  • 196696 03/20/07 01:48 C. Fernandez current snapshot
  • 196695 03/20/07 01:48 C. Fernandez Hall A general tools snapshot
  • 196694 03/20/07 01:45 C. Fernandez spot++ snapshot
  • 196691 03/20/07 01:39 Cornejo, J.C Raster current run 1757
  • 196690 03/20/07 01:38 Cornejo, J.C Raster current run 1755 (Pb)
  • 196682 03/20/07 00:03 Mina Katrama Shift Summary
  • Monday

  • 196678 03/19/07 23:40 Makis P Hall A Tools with corrected Left angle
  • 196676 03/19/07 23:39 Makis P Right angle is 21.440 (after camera fixing)
  • 196675 03/19/07 23:34 Makis P Hall access to fix Right arm S2 ADCs
  • 196670 03/19/07 22:37 Mina Kat IOC monitoring
  • 196669 03/19/07 22:25 POTOKAR Deadtime
  • 196666 03/19/07 22:06 Mina Kat spot++
  • 196665 03/19/07 22:01 Makis P target screen 1
  • 196657 03/19/07 21:09 Mina Kat Right spectrometer ROC1 crash
  • 196656 03/19/07 21:03 Mina Kat New monitor for left arm angle
  • 196654 03/19/07 20:54 Mina Kat coda crash- hall access- port server reboot
  • 196653 03/19/07 20:47 A. Camsonne run 1742 problem with ROC1
  • 196652 03/19/07 20:47 Makis P Run 1742 is NO Good/CODA problem
  • 196659 03/19/07 20:31 A. Camsonne Fastbus sequencer in bad state after first power cycling now is fine
  • 196658 03/19/07 20:11 R. Michaels Invalid data coincidence
  • 196655 03/19/07 19:59 R. Michaels alarms to ignore RMS=0
  • 196649 03/19/07 19:52 A. Camsonne ROC1 and right portserver crashed
  • 196648 03/19/07 19:11 Makis P Left arm angle/small discrepancy
  • 196645 03/19/07 18:57 Makis P Right arm angle/small discrepancy
  • 196644 03/19/07 18:42 A. Camsonne Getting all variables from 129.57.255.12
  • 196641 03/19/07 18:19 Mina Kat Spot++ for run 1738
  • 196638 03/19/07 17:58 Mina Kat IOC monitoring and "invalid channel name" messages
  • 196637 03/19/07 17:45 Mina Kat raster plots for run 1736
  • 196630 03/19/07 17:07 Mina Kat Spot++ for Pb+C target
  • 196627 03/19/07 16:54 Makis P target to Pb/start with 10uA
  • 196625 03/19/07 16:41 Mina Kat flatlines strip after IOC reboot
  • 196624 03/19/07 16:38 Mina Kat spot++ Carbon
  • 196635 03/19/07 16:32 R. Michaels alarms to ignore "Invalid" data
  • 196618 03/19/07 16:06 LeRose Rebooting right IOC's
  • 196617 03/19/07 15:58 LeRose Getting better
  • 196616 03/19/07 15:57 Shahinyan Target
  • 196613 03/19/07 15:45 LeRose Older Carbon
  • 196612 03/19/07 15:37 Vignote raster current and raster at target run 1731
  • 196611 03/19/07 15:35 LeRose Bad Steering?
  • 196623 03/19/07 15:33 R. Michaels voice alarms (again)
  • 196621 03/19/07 15:24 LeRose Shift Summary (16:24)
  • 196609 03/19/07 15:20 LeRose Time Test
  • 196606 03/19/07 15:16 Adaq test
  • 196608 03/19/07 14:20 R. Michaels test again
  • 196607 03/19/07 14:19 LeRose Time Test
  • 196602 03/19/07 14:09 LeRose Sigma y now really ok. 15:09
  • 196601 03/19/07 13:54 LeRose Still working on sigma_y! 14:51
  • 196600 03/19/07 13:46 LeRose Harp Scan y OK at 14:45
  • 196599 03/19/07 12:55 LeRose Harp Scan y still working
  • 196594 03/19/07 12:54 Shahinyan Target
  • 196598 03/19/07 12:48 LeRose Harp Scan y OK
  • 196590 03/19/07 12:34 Kalyan Right arm D1 current vs run number plots
  • 196587 03/19/07 12:21 Vignote raster current and raster at target run 1727
  • 196597 03/19/07 12:11 LeRose Time problems
  • 196586 03/19/07 12:06 Vignote raster current and raster at target run 1726
  • 196596 03/19/07 12:05 LeRose Harp Scan y too large
  • 196595 03/19/07 12:03 LeRose Restart Coda
  • 196581 03/19/07 11:23 LeRose Hall A Tools/beam
  • 196580 03/19/07 11:22 LeRose IOC monitoring
  • 196579 03/19/07 11:17 Vignote raster current and raster at target run 1725
  • 196576 03/19/07 10:54 Vignote raster current and raster at target run 1724
  • 196575 03/19/07 10:43 LeRose Raster
  • 196572 03/19/07 10:36 Vignote raster current and raster at target run 1723
  • 196571 03/19/07 10:15 LeRose Raster X vs Y (current)
  • 196565 03/19/07 08:49 LeRose Beam Back
  • 196562 03/19/07 08:34 LeRose Q1 Left Trip
  • 196560 03/19/07 08:30 Vignote raster current and raster at target run 1720
  • 196559 03/19/07 08:12 Shahinyan Target Status.
  • 196564 03/19/07 08:06 M. Stevens TECH Left Q1 trip, odd reset
  • 196557 03/19/07 08:06 C. Fernandez Owl Shift Summary
  • 196554 03/19/07 07:57 C. Fernandez Hall A general tools snapshot
  • 196563 03/19/07 07:55 LeRose No beam 09:00- 10:00
  • 196550 03/19/07 07:39 C. Fernandez Harp Scan
  • 196548 03/19/07 06:04 maria martin raster current and raster at target run 1716
  • 196543 03/19/07 05:34 maria martin raster current and raster at target run 1715
  • 196540 03/19/07 05:11 maria martin raster current and raster at target run 1714
  • 196537 03/19/07 04:48 C. Fernandez spot++ snapshot
  • 196536 03/19/07 04:47 C. Fernandez strip snapshot
  • 196535 03/19/07 04:46 C. Fernandez energy spread snapshot
  • 196534 03/19/07 04:46 C. Fernandez Current snapshot
  • 196533 03/19/07 04:41 maria martin raster current and raster at target run 1713
  • 196532 03/19/07 04:38 C. Fernandez Hall A general tools snapshot
  • 196529 03/19/07 04:18 maria martin raster current and raster at target run 1712
  • 196526 03/19/07 03:51 maria martin raster current and raster at target run 1711
  • 196523 03/19/07 03:30 maria martin raster current and raster at target run 1710
  • 196522 03/19/07 03:22 C. Fernandez spot++ snapshot
  • 196519 03/19/07 03:03 maria martin raster current and raster at target run 1709
  • 196517 03/19/07 02:45 atrig Trigger-Download
  • 196515 03/19/07 02:37 maria martin raster current and raster at target run 1708
  • 196512 03/19/07 02:11 maria martin raster current and raster at target run 1707
  • 196509 03/19/07 01:44 maria martin raster current and raster at target run 1706
  • 196506 03/19/07 01:21 maria martin raster current and raster at target run 1705
  • 196505 03/19/07 01:19 C. Fernandez Adaql8 rebooted
  • 196502 03/19/07 01:02 maria martin raster current and raster at target run 1704
  • 196501 03/19/07 00:49 C. Fernandez spot++ snapshot
  • 196498 03/19/07 00:40 maria martin raster current and raster at target run 1703
  • 196497 03/19/07 00:31 C. Fernandez energy spread snapshot (this time with the figure!)
  • 196496 03/19/07 00:30 C. Fernandez Energy spread snapshot
  • 196495 03/19/07 00:30 C. Fernandez strip snapshot
  • 196494 03/19/07 00:18 C. Fernandez Beam current snapshot
  • 196492 03/19/07 00:18 C. Fernandez Hall A general tools snapshot
  • 196489 03/19/07 00:01 bogdanw shift summary
  • Sunday

  • 196484 03/18/07 23:09 adaq HV screens
  • 196483 03/18/07 23:08 adaq tools for the run 1700
  • 196482 03/18/07 23:02 Milan Potoka spot++ run 1700
  • 196481 03/18/07 23:00 Milan Potoka spot++
  • 196479 03/18/07 22:47 Solvigno Move target to Pb
  • 196478 03/18/07 21:19 Solvigno Target to Carbon for harp scan
  • 196476 03/18/07 20:51 adaq raster for run 1699
  • 196474 03/18/07 20:42 solvigno Target status
  • 196473 03/18/07 20:40 solvigno Target chart (Pb in)
  • 196467 03/18/07 20:11 Solvigno Target to Pb
  • 196477 03/18/07 20:08 R. Michaels voice alarms turned off for the night
  • 196462 03/18/07 18:43 atrig Trigger-Download
  • 196457 03/18/07 17:13 Adaq unstable beam
  • 196460 03/18/07 17:04 M Jones Previous night's drift in EMISS caused by drifting right arm spec dipole current
  • 196454 03/18/07 16:25 E. Cisbani Emiss and right magnet field drifting, frozen ioc
  • 196452 03/18/07 16:05 Vignote day shift summary
  • 196449 03/18/07 15:49 Reinhold Target status (cont.)
  • 196448 03/18/07 15:48 Reinhold Target status
  • 196443 03/18/07 14:50 vignote spot ++ snapshot
  • 196445 03/18/07 14:24 R. Michaels alarms fixed, running on hapc5 now
  • 196440 03/18/07 11:56 Reinhold Thoughts on harp scans
  • 196437 03/18/07 11:19 maria martin new harp scan results
  • 196436 03/18/07 10:38 maria martin harp scan
  • 196435 03/18/07 09:57 Reinhold Target: moved to carbon (solid 2)
  • 196433 03/18/07 09:36 Reinhold Harp scan results
  • 196432 03/18/07 09:12 Reinhold Target: moved to empty target
  • 196430 03/18/07 09:05 Reinhold Target: T3 at 10 uA
  • 196427 03/18/07 08:57 Vignote spot++ snapshot
  • 196426 03/18/07 08:56 Vignote Hall A General Tools Snapshot
  • 196425 03/18/07 08:55 Vignote strip snapshot
  • 196424 03/18/07 08:54 Vignote energy spread
  • 196434 03/18/07 08:53 R. Michaels alarms not working on hapc7 (for now)
  • 196423 03/18/07 08:51 Reinhold Target: T3 at 20 uA
  • 196420 03/18/07 08:25 Reinhold Target status (cont.)
  • 196419 03/18/07 08:23 Reinhold Target status
  • 196416 03/18/07 08:09 C. Fernandez Owl Shift Summary
  • 196412 03/18/07 07:34 C. Fernandez spot++ snapshot
  • 196410 03/18/07 07:24 C. Fernandez energy spread snapshot
  • 196409 03/18/07 07:23 C. Fernandez strip snapshot
  • 196407 03/18/07 07:19 C. Fernandez Current snapshot
  • 196406 03/18/07 07:19 C. Fernandez Hall general tool snapshot
  • 196402 03/18/07 04:56 C. Fernandez Frozen Right Spectrometer: Rebooted
  • 196401 03/18/07 04:44 Cornejo run 1680 possibly junk
  • 196399 03/18/07 04:37 C. Fernandez spot++ snapshot
  • 196396 03/18/07 04:06 C. Fernandez energy spread snapshot
  • 196395 03/18/07 04:06 C. Fernandez strip snapshot
  • 196394 03/18/07 04:05 C. Fernandez current snapshot
  • 196393 03/18/07 04:05 C. Fernandez Hall A general tools snapshot
  • 196386 03/18/07 03:06 C. Fernandez spot++ snapshot
  • 196384 03/18/07 02:45 C. Fernandez 40 uA current
  • 196380 03/18/07 02:15 C. Fernandez Hall A current limitation: MCC Entry
  • 196379 03/18/07 02:11 C. Fernandez energy spread
  • 196378 03/18/07 02:10 C. Fernandez strip snapshot
  • 196377 03/18/07 02:10 C. Fernandez Current snapshot
  • 196376 03/18/07 02:09 C. Fernandez HAll A general tool snapshot
  • 196372 03/18/07 01:32 C. Fernandez Change to 20 uA
  • 196370 03/18/07 01:14 C. Fernandez spot++ snapshot
  • 196363 03/18/07 00:02 Solvigno Swing shift summary
  • Saturday

  • 196359 03/17/07 23:23 Solvigno Hall A general tools
  • 196358 03/17/07 23:22 Solvigno Energy spread snapshot
  • 196357 03/17/07 23:19 Milan Potoka spot++
  • 196354 03/17/07 23:16 S.Abrahamyan Target status
  • 196335 03/17/07 19:25 Solvigno Spot++
  • 196334 03/17/07 19:21 S.Abrahamyan Target status
  • 196333 03/17/07 19:20 Solvigno Energy spread screenshot
  • 196315 03/17/07 15:58 Vignote day shift summary
  • 196308 03/17/07 15:26 X. Zhan Target Vacuum Spike
  • 196307 03/17/07 15:22 X. Zhan Target: Pb3
  • 196303 03/17/07 14:34 vignote strip snapshot
  • 196302 03/17/07 14:33 Vignote energy spread
  • 196301 03/17/07 14:31 Adaq spot++
  • 196310 03/17/07 14:30 M Jones harp scan ( elog 1364671)
  • 196297 03/17/07 13:57 X. Zhan Target: Carbon
  • 196296 03/17/07 13:47 Vignote change to kin03
  • 196295 03/17/07 13:30 G.M. Urciuol BPM pedestal run
  • 196290 03/17/07 13:01 X. Zhan Target: Empty
  • 196289 03/17/07 11:59 Jack 24V raster supply replaced
  • 196271 03/17/07 09:34 Vignote spot++
  • 196266 03/17/07 09:01 vignote strip snapshot
  • 196265 03/17/07 09:00 vignote energy spread
  • 196264 03/17/07 08:52 vignote beam current
  • 196263 03/17/07 08:47 vignote hall A general tools
  • 196260 03/17/07 08:34 X. Zhan Target screen2
  • 196259 03/17/07 08:33 X. Zhan Target Screen
  • 196256 03/17/07 08:05 C. Fernandez Owl Shift Summary
  • 196251 03/17/07 07:46 C. Fernandez spot++ snapshot
  • 196248 03/17/07 07:16 C. Fernandez spot++ snapshot
  • 196254 03/17/07 06:55 M Jones Fixing raster power supply
  • 196245 03/17/07 06:51 C. Fernandez Energy spread snapshot
  • 196244 03/17/07 06:51 C. Fernandez strip snapshot
  • 196243 03/17/07 06:50 C. Fernandez Beam current snapshot
  • 196242 03/17/07 06:50 C. Fernandez Hall A general tools snapshot
  • 196241 03/17/07 06:49 R. Michaels Pb target is fine
  • 196240 03/17/07 06:47 C. Fernandez spot++ snapshot, 40 uA
  • 196237 03/17/07 06:39 C. Fernandez CORRECTION: spot snapshot, 20 uA
  • 196236 03/17/07 06:33 C. Fernandez spot++ snapshot, 20 uA
  • 196235 03/17/07 06:29 C. Fernandez spot++ snapshot
  • 196234 03/17/07 06:27 C. Fernandez Beam on, production
  • 196232 03/17/07 06:18 R. Michaels move to Pb#3
  • 196230 03/17/07 06:17 R. Michaels spot++ check
  • 196228 03/17/07 06:13 R. Michaels Raster repair
  • 196227 03/17/07 05:21 C. Fernandez Controlled access granted
  • 196226 03/17/07 05:06 C. Fernandez Raster died. Asked for control access
  • 196225 03/17/07 04:59 R. Michaels raster failure, move to empty target
  • 196221 03/17/07 04:43 C. Fernandez spot++ snapshot
  • 196214 03/17/07 02:16 J Lopez Emiss - KIN 5 - Looks fine
  • 196206 03/17/07 01:58 C. Fernandez spot++ snapshot
  • 196205 03/17/07 01:56 C. Fernandez strip snapshot
  • 196204 03/17/07 01:56 C. Fernandez Energy spread snapshot
  • 196203 03/17/07 01:55 C. Fernandez Beam Current Snapshot
  • 196202 03/17/07 01:54 C. Fernandez Hall A General Tool Snapshot
  • 196211 03/17/07 01:49 J Lopez Emiss spectrum (Pb+C) KIN 7
  • 196197 03/17/07 01:22 C. Fernandez spot++ snapshot
  • 196194 03/17/07 01:03 R. Michaels re: R-arm IOC freeze on swing
  • 196193 03/17/07 01:01 Cornejo PS3 for run 1613 1614 are wrong
  • 196185 03/17/07 00:05 weinstein shift summary
  • Friday

  • 196182 03/16/07 23:51 Shahinyan Target
  • 196181 03/16/07 23:39 weinstein spot++
  • 196179 03/16/07 23:04 weinstein better harp scan
  • 196178 03/16/07 22:52 weinstein beam size
  • 196177 03/16/07 22:48 weinstein harp scan
  • 196176 03/16/07 22:36 Shahinyan Target move
  • 196174 03/16/07 22:17 weinstein general tools
  • 196173 03/16/07 22:16 weinstein srtip charts
  • 196160 03/16/07 19:45 weinstein hall a general tools gui
  • 196159 03/16/07 19:43 weinstein Right arm ioc reboot
  • 196157 03/16/07 19:14 weinstein Right arm ioc reboot
  • 196154 03/16/07 18:57 weinstein new voice alarms
  • 196152 03/16/07 18:37 Aniol returned raster to 4mm x 6mm
  • 196147 03/16/07 17:41 weinstein general tools
  • 196146 03/16/07 17:40 weinstein strip charts
  • 196134 03/16/07 16:20 weinstein spot++
  • 196132 03/16/07 16:18 Shahinyan Target status
  • 196129 03/16/07 16:05 bogdan shift summary
  • 196124 03/16/07 15:24 X. Zhan Spot++ run 1591
  • 196127 03/16/07 14:46 R. Michaels how to see if hitting frame with spot++
  • 196118 03/16/07 13:45 X. Zhan Target moved to Lead
  • 196116 03/16/07 13:24 X. Zhan Beam monitor screen
  • 196115 03/16/07 13:21 X. Zhan Spot++ run 1588
  • 196111 03/16/07 11:44 Ole Hansen Note that runs 1553-1555 are labeled as junk
  • 196109 03/16/07 10:02 X. Zhan Target: Carbon
  • 196108 03/16/07 09:50 Maria Martin Raster at target run 1586
  • 196107 03/16/07 09:34 Maria Martin Replay run 1586
  • 196110 03/16/07 09:30 R. Michaels new voice alarms
  • 196106 03/16/07 09:00 X. Zhan Spot++ run 1585
  • 196105 03/16/07 08:56 X. Zhan Target Screen2
  • 196104 03/16/07 08:55 X. Zhan Target Screen
  • 196103 03/16/07 08:54 X. Zhan HV Right Arm Screen
  • 196102 03/16/07 08:52 X. Zhan HV Left Screen
  • 196101 03/16/07 08:49 X. Zhan Beamtool screen
  • 196100 03/16/07 08:47 X. Zhan Hall A General Tools Screen
  • 196099 03/16/07 08:45 Xiaohui Spot++
  • 196097 03/16/07 08:05 C. Fernandez Owl Shift Summary
  • 196083 03/16/07 05:23 C. Fernandez Energy spread strip
  • 196080 03/16/07 04:56 C. Fernandez strip snapshot
  • 196079 03/16/07 04:55 C. Fernandez 40 uA current beam snapshot
  • 196078 03/16/07 04:55 C. Fernandez Hall A general tools snapshot
  • 196077 03/16/07 04:54 C. Fernandez spot++ snapshot
  • 196072 03/16/07 03:23 J Lopez Raster at target in Carbon (runs 1553-1561)
  • 196059 03/16/07 01:38 C. Fernandez Spot++
  • 196056 03/16/07 01:15 C. Fernandez Beam current snapshot
  • 196055 03/16/07 01:12 C. Fernandez Strip Snapshot
  • 196054 03/16/07 01:11 C. Fernandez Hall A General Tool Snapshot
  • Thursday

  • 196045 03/15/07 23:59 jpchen swing shift summary
  • 196043 03/15/07 23:39 jpchen general tools and beam strip chart
  • 196038 03/15/07 22:53 jpchen lead flow strip chart
  • 196037 03/15/07 22:48 jpchen run 1563 is 20 uA on Pb at kin07
  • 196036 03/15/07 22:48 jpchen run 1562 is 10 uA on Pb at kin 07.
  • 196030 03/15/07 22:08 Ole Hansen Target to lead
  • 196028 03/15/07 21:57 jpchen strip chart and general tools
  • 196022 03/15/07 21:11 Ole Hansen Target to carbon
  • 196021 03/15/07 19:26 Ole Hansen Target to empty
  • 196018 03/15/07 18:50 Ole Hansen Target vacuum excursion
  • 196017 03/15/07 18:11 jpchen beam energy spread was too high
  • 196012 03/15/07 16:09 Higinbotham Shift Summary
  • 196011 03/15/07 16:03 Shahinyan Target status
  • 196008 03/15/07 15:14 Higinbotham Target Vacuum Spike
  • 196007 03/15/07 14:57 Higinbotham Tools Screen
  • 196005 03/15/07 14:56 Higinbotham spot ++
  • 196003 03/15/07 14:55 Higinbotham spot++
  • 195997 03/15/07 13:57 Shahinyan target move
  • 196006 03/15/07 13:56 M Jones Harp results ( from elog 1364348)
  • 196001 03/15/07 13:48 M Jones Another check for spectrometer IOC problems
  • 195996 03/15/07 09:28 LeRose Rebooting Spectrometer IOC's Procedure
  • 195995 03/15/07 08:55 Higinbotham Left HRS Problem Started at 4am
  • 195994 03/15/07 08:12 Chiranjib target moved to empty
  • 195993 03/15/07 08:06 Cesar Fernan Owl Shift Summary
  • 195990 03/15/07 07:48 Cesar Fenand Strip tool graph
  • 195989 03/15/07 07:48 Cesar Fernan General Tool
  • 195967 03/15/07 02:29 Chiranjib target moved to Pb
  • 195964 03/15/07 01:36 Chiranjib target moved to C
  • 195963 03/15/07 01:30 Cesar Fernan Harp Scan
  • 195962 03/15/07 01:22 Chiranjib target moved to empty
  • 195961 03/15/07 01:12 Cesar Fernan Beam turned off
  • 195959 03/15/07 01:01 Cesar Fernan Beam current reduced
  • 195956 03/15/07 00:55 Cesar Fernan Beam Current Reduced
  • 195951 03/15/07 00:02 A. Puckett swing shift summary
  • Wednesday

  • 195947 03/14/07 23:47 A. Puckett Hall A Tools and BCM snapshots
  • 195946 03/14/07 23:45 A. Puckett SLI stripchart grab
  • 195945 03/14/07 23:34 A. Puckett SLI screengrab
  • 195940 03/14/07 22:54 A. Puckett spot check
  • 195923 03/14/07 19:24 A. Puckett screen shots from HARP scan
  • 195922 03/14/07 19:13 A. Puckett Right arm angle for run #'s 1521, 1522
  • 195917 03/14/07 18:22 A. Puckett SLI/BCM stripcharts
  • 195914 03/14/07 18:04 A. Puckett Changed to kin. 07
  • 195909 03/14/07 17:29 A. Puckett spot check with 10 muA on lead
  • 195903 03/14/07 16:31 A. Puckett SLI monitor screengrab
  • 195907 03/14/07 16:22 M Jones Beam will be down 3/15 9am-noon for Hall B pass change
  • 195900 03/14/07 15:01 LeRose Shift Summary
  • 195894 03/14/07 14:22 LeRose Rebooting Spectrometer IOC's Procedure
  • 195893 03/14/07 14:10 LeRose Beam Energy Shift
  • 195883 03/14/07 13:18 LeRose Beam Energy Shifted
  • 195882 03/14/07 13:15 LeRose Kin 01
  • 195880 03/14/07 12:58 atrig Trigger-Download
  • 195886 03/14/07 12:52 LeRose Quads Cycled
  • 195878 03/14/07 11:42 G.M. Urciuol The Beam hits the frame
  • 195866 03/14/07 09:32 R. Michaels re: spikes (hitting something)
  • 195871 03/14/07 09:18 LeRose Sigma y now ok
  • 195869 03/14/07 08:59 LeRose Harp Scan y too large
  • 195861 03/14/07 08:31 S. Sirca Spikes in raster pattern on carbon
  • 195865 03/14/07 08:27 LeRose 1C12 OTR removed
  • 195860 03/14/07 08:08 Eric Voutier Shit summary
  • 195858 03/14/07 07:56 Eric Voutier kin#12 settings
  • 195855 03/14/07 07:41 Eric Voutier NMR Lock
  • 195852 03/14/07 07:14 Eric Voutier Magnet settings / kin12
  • 195850 03/14/07 06:11 Eric Voutier Magnet problems
  • 195846 03/14/07 01:36 byung target 'C'
  • 195840 03/14/07 00:03 E.Chudakov Swing shift summary
  • Tuesday

  • 195837 03/13/07 23:56 A. Puckett target screen grabs 2
  • 195836 03/13/07 23:54 A. Puckett target screen grabs 1
  • 195835 03/13/07 23:53 A. Puckett T3 rate / Coulomb for lead runs, swing shift
  • 195832 03/13/07 23:16 E.Chudakov TOOLS, strip chart
  • 195829 03/13/07 22:38 A. Puckett Lead at 40 muA
  • 195827 03/13/07 22:25 A. Puckett Lead at 20 muA
  • 195825 03/13/07 22:14 A. Puckett Lead at 10 muA, run 1494
  • 195820 03/13/07 21:32 E.Chudakov HRSR reboot
  • 195815 03/13/07 20:55 E.Chudakov C12 energy spread
  • 195814 03/13/07 20:43 A. Puckett spot/raster++ after beam restoration
  • 195813 03/13/07 20:40 A. Puckett beam back, replay for Carbon, funny effect in raster histograms.
  • 195809 03/13/07 20:14 E.Chudakov SPOT for carbon target
  • 195806 03/13/07 19:57 E.Chudakov Harp scan results
  • 195805 03/13/07 18:08 J Lopez KIn6 looks fine
  • 195804 03/13/07 16:12 J Lopez Be careful with Lead target
  • 195802 03/13/07 16:03 jpchen day shift summary
  • 195800 03/13/07 14:42 R. Michaels ignore_err in scaler helicity
  • 195801 03/13/07 13:52 jose tech /general tools, spectrometers screen
  • 195795 03/13/07 12:26 G.M. Urciuol Problems with Right Arm and Missing energy shift
  • 195794 03/13/07 11:13 S. Sirca GenTools / Transition to kin #13
  • 195787 03/13/07 08:02 Jon Miller Shift Summary
  • 195786 03/13/07 07:35 Cornejo Raster at Pb target run 1483
  • 195783 03/13/07 07:26 Jon Miller General Tools + Strip Chart
  • 195782 03/13/07 07:25 Jon Miller SLI
  • 195777 03/13/07 06:44 Cornejo Raster on Pb target 1481
  • 195775 03/13/07 06:15 Cornejo Target vacumm alarm
  • 195771 03/13/07 05:33 Cornejo Raster on Pb target run 1478
  • 195766 03/13/07 04:46 Jon Miller Spot++
  • 195765 03/13/07 04:21 Jon Miller General Tools + Strip Chart
  • 195758 03/13/07 02:45 Cornejo Raster at Pb target run 1473
  • 195751 03/13/07 01:02 Jon Miller Spot++
  • 195750 03/13/07 01:01 Jon Miller General Tools + Strip Chart
  • 195747 03/13/07 00:54 Jonathan Mil Alarm
  • 195743 03/13/07 00:02 weinstein shift summary
  • Monday

  • 195741 03/12/07 23:44 A. Puckett T3 Rates by run, swing shift
  • 195745 03/12/07 23:36 J LOPEZ Sunday logs are gone
  • 195738 03/12/07 23:29 weinstein spot++
  • 195735 03/12/07 22:56 weinstein strip chart
  • 195728 03/12/07 21:05 weinstein general tools gui
  • 195724 03/12/07 18:59 A. Puckett target screen grabs, 2
  • 195723 03/12/07 18:57 A. Puckett target screen grabs, 1
  • 195722 03/12/07 18:47 weinstein strip charts
  • 195716 03/12/07 16:44 R. Michaels coinc time (another view)
  • 195713 03/12/07 16:16 weinstein left arm IOC reboot
  • 195712 03/12/07 16:11 R. Michaels L-arm TC messed up, fix
  • 195710 03/12/07 16:07 jpchen Endrun 1458, no content.
  • 195709 03/12/07 16:05 jpchen Day Shift Summary
  • 195708 03/12/07 16:02 T. Ngo New trigger
  • 195705 03/12/07 16:01 atrig Trigger-Download
  • 195702 03/12/07 15:10 jpchen beam SLI strip chart
  • 195701 03/12/07 15:08 jpchen Hall A general tools screen
  • 195698 03/12/07 14:35 S. Sirca Trigger reload, coda restart
  • 195696 03/12/07 14:31 atrig Trigger-Download
  • 195694 03/12/07 14:22 R. Michaels Coinc. time check
  • 195683 03/12/07 11:28 jpchen end of run 1449/1450
  • 195677 03/12/07 09:46 R. Michaels Pb target #3 looking good
  • 195675 03/12/07 09:17 jpchen beam strip chart
  • 195673 03/12/07 09:14 jpchen reboot ioc for right magnets and NMR.
  • 195669 03/12/07 08:25 Chiranjib Vacuum spike
  • 195666 03/12/07 08:02 Jon Miller Shift Summary
  • 195660 03/12/07 06:54 Jon Miller SLI
  • 195659 03/12/07 06:53 Jon Miller Spot++
  • 195654 03/12/07 05:31 Jon Miller General Tools + Strip Chart
  • 195640 03/12/07 02:42 Cornejo Target status screen shots
  • 195639 03/12/07 02:41 Cornejo Pb Target Screen Shots
  • 195631 03/12/07 00:54 Jon Miller General Tools + Strip Chart
  • 195625 03/12/07 00:02 bogdanw shift summary
  • Sunday

  • 195623 03/11/07 23:47 adaq SLI
  • 195621 03/11/07 23:27 Mina Kat spot++ 2nd try
  • 195620 03/11/07 23:25 Mina Kat spot++
  • 195617 03/11/07 23:13 Adaq tools
  • 195616 03/11/07 23:12 adaq beam screen
  • 195611 03/11/07 21:44 adaq beam screen
  • 195606 03/11/07 20:49 adaq spot++ for Bi
  • 195602 03/11/07 20:31 Makis P Moved tgt to Bi
  • 195600 03/11/07 20:26 adaq SLI
  • 195598 03/11/07 20:13 Makis P Target charts
  • 195596 03/11/07 20:10 Makis P Target screens
  • 195595 03/11/07 20:03 Mina Kat Peculiar looking spot++
  • 195594 03/11/07 19:34 adaq beam screen
  • 195593 03/11/07 19:34 adaq tools kin 6
  • 195591 03/11/07 19:29 Makis Moved target to Carbon
  • 195590 03/11/07 19:15 Makis P Target to EMPTY/Harp Scan
  • 195586 03/11/07 18:17 Mina Kat spot++
  • 195579 03/11/07 16:53 adaq HV screens
  • 195578 03/11/07 16:51 adaq tools
  • 195577 03/11/07 16:49 Mina Kat spot++
  • 195572 03/11/07 16:05 S. Sirca Day shift summary
  • 195554 03/11/07 12:19 S. Sirca Postponing harp scan at target
  • 195545 03/11/07 10:37 S. Sirca Annoying IOC freezes for HRSR
  • 195542 03/11/07 10:09 S. Sirca BCM snapshot
  • 195538 03/11/07 09:40 S. Sirca General Tools snapshot
  • 195536 03/11/07 09:25 S. Sirca SLI snapshot
  • 195535 03/11/07 09:16 S. Sirca Raster size run 1397
  • 195530 03/11/07 08:30 S. Sirca BTA confused by DST
  • 195527 03/11/07 08:05 Markowitz shift summary
  • 195503 03/11/07 01:55 Adaq strip chart
  • 195502 03/11/07 01:54 Markowitz Hall A tools, bcm screens
  • 195490 03/11/07 00:04 bogdanw shift summary
  • Saturday

  • 195485 03/10/07 23:08 adaq beam screen
  • 195484 03/10/07 23:06 adaq HV screens
  • 195483 03/10/07 23:05 Makis P spot++ at start of I-scan Pb
  • 195482 03/10/07 23:04 bogdanw tools
  • 195480 03/10/07 22:54 Mina Katrama Moved target to Pb#3
  • 195479 03/10/07 22:33 Mina Katrama Move target to empty
  • 195473 03/10/07 21:30 adaq He level
  • 195461 03/10/07 20:11 Makis Petrat spot++ (run 1368 8pm)
  • 195457 03/10/07 19:43 Adaq beam screen
  • 195456 03/10/07 19:40 adaq raster run 1366
  • 195453 03/10/07 19:20 Mina Katrama target charts (continued)
  • 195452 03/10/07 19:18 Mina Katrama Target screens
  • 195450 03/10/07 19:04 bogdanw tools
  • 195447 03/10/07 18:26 Mina Katrama Target moved to Bi
  • 195446 03/10/07 18:22 Mina Katr spott++
  • 195445 03/10/07 18:18 mina Katrama move to empty position for harp scan
  • 195438 03/10/07 16:04 S. Sirca Day shift summary
  • 195433 03/10/07 15:25 S. Sirca Harp scan at target
  • 195429 03/10/07 14:33 X. Zheng target to Bi
  • 195428 03/10/07 13:55 X. Zheng target to empty
  • 195424 03/10/07 13:26 S. Sirca File server down
  • 195421 03/10/07 13:02 S. Sirca Summary of I-scan runs
  • 195415 03/10/07 12:05 S. Sirca SLI snapshot and warning
  • 195413 03/10/07 12:03 X. Zheng raster current pattern for latest run
  • 195411 03/10/07 11:42 X. Zheng Target to Bi now
  • 195409 03/10/07 10:39 S. Sirca General Tools snapshot
  • 195408 03/10/07 10:29 S. Sirca Poor beam quality for I-scan runs
  • 195402 03/10/07 08:58 X. Zheng Target move back to Bi
  • 195401 03/10/07 08:58 S. Sirca New MCC values for raster
  • 195396 03/10/07 08:44 S. Sirca Raster pattern on 12C
  • 195394 03/10/07 08:34 S. Sirca Reverting to 12C target
  • 195391 03/10/07 07:59 Markowitz shift summary
  • 195390 03/10/07 07:59 Margaziotis Moved to Bismuth target
  • 195385 03/10/07 07:11 Markowitz tools and bcm screenshots
  • 195383 03/10/07 07:09 Adaq current strip chart
  • 195371 03/10/07 03:19 Markowitz raster on
  • 195369 03/10/07 03:04 markowitz raster off
  • 195367 03/10/07 02:53 Markowitz MCC calibrating tune lock
  • 195365 03/10/07 02:47 Markowitz move to kin 4
  • 195364 03/10/07 02:45 Markowitz carbon target in
  • 195363 03/10/07 02:33 jose Always raster on with Pb and Bi targets
  • 195362 03/10/07 02:14 J Lopez Target status from physics replay
  • 195361 03/10/07 02:08 Markowitz sli improved
  • 195360 03/10/07 02:08 Markowitz target back to empty for HARP scan
  • 195359 03/10/07 01:58 Markowitz target vacuum
  • 195358 03/10/07 01:56 Adaq Target status from physics replay
  • 195357 03/10/07 01:55 Adaq sli screenshot
  • 195356 03/10/07 01:52 Markowitz beam back
  • 195355 03/10/07 01:51 Markowitz spot++ screen
  • 195353 03/10/07 01:44 Markowitz target to position 4 (Pb)
  • 195352 03/10/07 01:30 JOAQUIN LOPE Target status from physics replay
  • 195351 03/10/07 00:10 ipom shift summary
  • Friday

  • 195349 03/09/07 21:25 ipom RHRS IOC Froze
  • 195348 03/09/07 20:22 ipom General tools screenshot
  • 195347 03/09/07 18:14 ipom RHRS IOC is frozen
  • 195346 03/09/07 17:38 Aniol recheck PB#4 at kin01
  • 195345 03/09/07 17:32 ipom Moving to Kin pt no. 1
  • 195343 03/09/07 16:52 R. Michaels Emiss of lead
  • 195342 03/09/07 16:07 ipom General tools screenshot
  • 195341 03/09/07 16:04 LeRose Shift Summary
  • 195339 03/09/07 15:49 LeRose Still no beam
  • 195338 03/09/07 15:19 R. Michaels Add HCBMW*
  • 195337 03/09/07 14:56 Elena Kh. Target screens
  • 195336 03/09/07 13:23 R. Michaels added HABMWCALC.J
  • 195335 03/09/07 12:42 R. Michaels X-Y from current scan runs.
  • 195332 03/09/07 12:28 R. Michaels poss. hole in lead target
  • 195331 03/09/07 12:21 R. Michaels access finished, beam status
  • 195329 03/09/07 12:16 LeRose No beam
  • 195327 03/09/07 11:44 R. Michaels L-arm ctime fixed
  • 195322 03/09/07 11:24 LeRose Beam Strip Tool
  • 195321 03/09/07 11:22 LeRose Tools/Beam current
  • 195319 03/09/07 11:04 LeRose Left arm controls frozen/rebooted
  • 195318 03/09/07 10:57 R. Michaels possible degradation of targets ?
  • 195317 03/09/07 10:41 LeRose Kin 4
  • 195316 03/09/07 10:30 R. Michaels pulser check of raster current
  • 195313 03/09/07 10:23 Elena Kh. Empty target
  • 195311 03/09/07 10:20 R. Michaels prescales, spin dance, access
  • 195310 03/09/07 10:20 LeRose Beam OFF/Controlled Access
  • 195308 03/09/07 09:59 R. Michaels pion rejector: nobody home
  • 195307 03/09/07 09:59 LeRose Half wave plate IN
  • 195306 03/09/07 09:57 R. Michaels Coinc time on L-arm absent
  • 195305 03/09/07 09:43 R. Michaels apparent Y non-uniform raster
  • 195303 03/09/07 09:11 R. Michaels adjust prescale factors
  • 195299 03/09/07 08:28 Aniol bismuth Iscan
  • 195298 03/09/07 08:23 LeRose Start runs labeled incorrectly
  • 195295 03/09/07 08:02 Markowitz shift summary
  • 195290 03/09/07 07:45 Markowitz target vacuum
  • 195289 03/09/07 07:34 Markowitz move to Bi target
  • 195276 03/09/07 06:10 Markowitz strip charts
  • 195261 03/09/07 04:33 adaq current screen shots
  • 195249 03/09/07 03:03 adaq strip chart
  • 195237 03/09/07 01:28 cusanno Raster current monitoring and situation
  • 195236 03/09/07 01:28 cusanno coincidence time in detector replay
  • 195230 03/09/07 00:24 markowitz target screenshots
  • 195229 03/09/07 00:23 Markowitz sli stripchart, tools
  • 195228 03/09/07 00:22 markowitz sli
  • Thursday

  • 195224 03/08/07 23:57 ipom Shift summary
  • 195221 03/08/07 23:50 Aniol spikes in raster return
  • 195213 03/08/07 23:28 ipom spot at (-4,4)
  • 195212 03/08/07 23:26 ipom spot on (-4,3)
  • 195209 03/08/07 23:23 ipom spot on (-4,1)
  • 195206 03/08/07 23:21 ipom hitting the frame in the y direction
  • 195201 03/08/07 23:01 ipom Changing to Pb
  • 195199 03/08/07 22:51 ipom start taking data
  • 195197 03/08/07 22:29 ipom SLI screenshot
  • 195196 03/08/07 22:29 ipom HARP Scan
  • 195195 03/08/07 22:20 ipom General tools screenshot
  • 195188 03/08/07 20:31 ipom Beam will not improve soon. starting runs.
  • 195187 03/08/07 20:12 ipom SLI Screenshot
  • 195186 03/08/07 20:11 ipom SLI Screenshot
  • 195185 03/08/07 20:01 ipom General tools screenshot
  • 195184 03/08/07 19:59 ipom No beam
  • 195183 03/08/07 18:56 ipom General tools screenshot
  • 195182 03/08/07 16:34 ipom SLI screenshot
  • 195180 03/08/07 16:15 Lindgren Shift Summary
  • 195179 03/08/07 15:32 Elena Kh. Empty target
  • 195178 03/08/07 15:08 Higinbotham SLI Picture
  • 195177 03/08/07 14:54 Higinbotham Hall A Tools Screen
  • 195148 03/08/07 12:03 R. Michaels flatness of raster current
  • 195146 03/08/07 11:41 Elena Kh. Target screens
  • 195145 03/08/07 11:10 Elena Kh. C target
  • 195144 03/08/07 10:58 R. Michaels raster readout adjustment -- finished
  • 195139 03/08/07 10:37 Adaq TECH and raster
  • 195127 03/08/07 09:56 Lindgren Raster repaired
  • 195125 03/08/07 08:20 Markowitz shift summary
  • 195124 03/08/07 05:27 Markowitz empty target in
  • 195122 03/08/07 05:08 Markowitz mcc takes beam for raster
  • 195121 03/08/07 05:00 Markowitz beam to 85 uA
  • 195120 03/08/07 04:36 markowitz beam current to 80 uA
  • 195115 03/08/07 03:56 Markowitz adjusted prescales
  • 195110 03/08/07 03:46 markowitz sli screen
  • 195105 03/08/07 03:38 Markowitz beam current, tools screens
  • 195103 03/08/07 03:35 Markowitz Pb target in
  • 195099 03/08/07 00:02 ipom shift summary
  • Wednesday

  • 195098 03/07/07 22:49 ipom General tools screenshot
  • 195097 03/07/07 21:34 ipom General tools screenshot
  • 195096 03/07/07 19:57 ipom General tools screenshot
  • 195095 03/07/07 18:30 ipom Right NMR power cycle
  • 195094 03/07/07 18:11 ipom IOC right HRS reboot
  • 195093 03/07/07 17:29 R. Michaels poss. solns to raster current spikes
  • 195092 03/07/07 17:27 ipom General tools screenshot
  • 195091 03/07/07 17:26 ipom SLI screenshot
  • 195090 03/07/07 17:24 atrig Trigger-Download
  • 195088 03/07/07 17:04 R. Michaels raster current pattern explained
  • 195085 03/07/07 16:52 R. Michaels check raster pedestals
  • 195081 03/07/07 16:04 ipom General tools screenshot
  • 195080 03/07/07 15:58 B. Craver day shift summary
  • 195074 03/07/07 14:53 R. Michaels raster checks
  • 195062 03/07/07 13:42 B. Craver Controlled Access
  • 195060 03/07/07 13:03 B. Craver General tools screenshot
  • 195059 03/07/07 13:02 B. Craver SLI sripchart
  • 195058 03/07/07 11:54 Elena Kh. Target screens
  • 195057 03/07/07 11:48 R. Michaels raster current on R-arm
  • 195054 03/07/07 10:37 Elena Kh. Empty target
  • 195053 03/07/07 10:35 B. Craver beam down
  • 195047 03/07/07 08:32 B. Craver (TECH) Ar pressure
  • 195044 03/07/07 08:04 B. Craver Argon pressure decreasing
  • 195043 03/07/07 08:02 Jon Miller Shift Summary
  • 195039 03/07/07 07:27 Jon Miller Hall A tools + Strip Chart
  • 195034 03/07/07 06:46 Jon Miller Spot++
  • 195033 03/07/07 06:36 Jon Miller SLI - Energy Spread
  • 195026 03/07/07 04:16 Jon Miller General Tools
  • 195025 03/07/07 04:15 byung target snapshot
  • 195024 03/07/07 04:06 byung target 'C'
  • 195015 03/07/07 00:57 Jon Miller Beam Size
  • 195012 03/07/07 00:36 Jon Miller General Tools
  • 195008 03/07/07 00:00 ipom shift summary
  • Tuesday

  • 195007 03/06/07 23:45 B. Craver spot++
  • 195006 03/06/07 23:42 ipom start production
  • 195004 03/06/07 23:33 B. Craver Target to Pb position
  • 195003 03/06/07 23:24 ipom General tools screenshot
  • 195002 03/06/07 23:17 ipom beam width status
  • 195001 03/06/07 23:13 B. Craver Target screens
  • 195000 03/06/07 22:57 B. Craver harp scan error
  • 194992 03/06/07 21:59 ipom General tools screenshot
  • 194991 03/06/07 21:41 ipom HARP scan
  • 194989 03/06/07 21:32 ipom Beam width problem
  • 194986 03/06/07 20:53 ipom General tools screenshot
  • 194985 03/06/07 19:15 ipom HV off on beamline channels
  • 194984 03/06/07 19:04 ipom Beam width problem
  • 194983 03/06/07 17:51 ipom vertical beam width problem
  • 194981 03/06/07 17:11 ipom HARP scan
  • 194977 03/06/07 16:52 B. Craver Target to empty
  • 194976 03/06/07 16:03 jpchen beam strip chart
  • 194974 03/06/07 16:00 jpchen day shift summary
  • 194966 03/06/07 13:58 jpchen beam tool strip chart
  • 194965 03/06/07 13:42 weinstein/jp spectrometer angles changed
  • 194962 03/06/07 13:36 Elena Kh. ERS_flow>12 g/s
  • 194959 03/06/07 13:00 weinstein strip chart
  • 194952 03/06/07 12:08 Elena Kh. Target screens
  • 194947 03/06/07 11:36 R. Michaels re: raster current measurements
  • 194946 03/06/07 11:30 jpchen started beamtool1 strip chart
  • 194943 03/06/07 11:07 jpchen left ioc reboot, control computer frozen
  • 194933 03/06/07 09:57 R. Michaels R-arm bottom HV, and hatsv3
  • 194930 03/06/07 09:26 R. Michaels halla_scaler_process.tcl fix
  • 194929 03/06/07 09:14 jpchen raster check with spot++
  • 194925 03/06/07 08:01 Jon Miller Shift Summary
  • 194922 03/06/07 07:12 Jon Miller General Tools
  • 194921 03/06/07 07:05 JonMiller SLI
  • 194918 03/06/07 06:09 Jon Miller Spot++
  • 194915 03/06/07 04:47 byung Target 'Pb'
  • 194914 03/06/07 04:46 byung target snapshot
  • 194912 03/06/07 04:22 Jon Miller Snapshots
  • 194907 03/06/07 01:52 Jon Miller SLI
  • 194906 03/06/07 01:46 Jon Miller Moved Spectrometer
  • 194904 03/06/07 00:39 byung target - C
  • 194903 03/06/07 00:38 byung target snapshot
  • 194902 03/06/07 00:29 Jon Miller Snapshots
  • 194898 03/06/07 00:00 ipom swing shift summary
  • Monday

  • 194897 03/05/07 22:45 B. Craver beam spot
  • 194896 03/05/07 22:40 Aniol horizontal raster ok at (-4,2)
  • 194893 03/05/07 22:32 Adaq raster at (-3.5,2.)
  • 194891 03/05/07 22:26 B. Craver adjusting beam horizontally
  • 194889 03/05/07 22:23 B. Craver spot_R plots
  • 194887 03/05/07 22:19 B. Craver more raster plots
  • 194886 03/05/07 22:19 ipom general tools screenshot
  • 194880 03/05/07 20:33 ipom current is dropped to 47uA
  • 194879 03/05/07 19:44 B. Craver non-uniformity in horizontal raster
  • 194875 03/05/07 19:27 B. Craver strange spot pattern
  • 194869 03/05/07 19:07 ipom pb led raster iscan
  • 194867 03/05/07 19:00 ipom sli tools screenshot
  • 194858 03/05/07 18:35 Ipom start Iscan
  • 194846 03/05/07 16:01 jpchen day shift summary
  • 194845 03/05/07 15:41 jpchen left spectrometer set to 2.216 GeV/c
  • 194844 03/05/07 11:28 Elena Kh. Target screens
  • 194843 03/05/07 10:42 jpchen 12C I scan completed, beam off for Hall C tuning
  • 194836 03/05/07 10:09 jpchen SLI strip chart
  • 194825 03/05/07 09:44 jpchen sieve out field to 0%: 2.635
  • 194815 03/05/07 08:03 Jon Miller Shift Summary
  • 194809 03/05/07 07:01 Jon Miller Beam Data
  • 194803 03/05/07 05:48 Jon Miller SLI Control
  • 194800 03/05/07 05:28 Jon Miller General Tools
  • 194792 03/05/07 04:09 Jon Miller Raster Size
  • 194760 03/05/07 00:57 Jon Miller Beam Size
  • 194758 03/05/07 00:50 Jon Miller Raster Size
  • 194753 03/05/07 00:10 Solvigno Shift summary
  • Sunday

  • 194750 03/04/07 23:49 solvigno harp scan for current situation
  • 194747 03/04/07 23:34 X. Zhan Raster Size
  • 194738 03/04/07 23:12 solvigno snapshop of SLI and Hall A general tools
  • 194733 03/04/07 22:33 X. Zhan Target Screen2
  • 194731 03/04/07 22:32 X. Zhan Target Screen
  • 194730 03/04/07 22:32 solvigno Raster size at max.
  • 194725 03/04/07 22:15 X. Zhan Target: Carbon
  • 194724 03/04/07 22:11 Solvigno Beam spot size
  • 194723 03/04/07 21:12 Solvigno Beam status
  • 194722 03/04/07 19:16 atrig Trigger-Download
  • 194721 03/04/07 16:00 X. Zheng Shift summary
  • 194719 03/04/07 13:57 X. Zheng Note on target pressure
  • 194718 03/04/07 13:22 B. Craver Target screen shot
  • 194717 03/04/07 13:22 X. Zheng Hall A General Tools Screen
  • 194716 03/04/07 12:07 jtewing polarity change-rt arm move
  • 194715 03/04/07 10:32 B. Craver target to empty
  • 194714 03/04/07 10:10 X. Zheng Hall in controlled access
  • 194713 03/04/07 10:08 B. Craver target pressure within limits
  • 194712 03/04/07 09:53 B. Craver target pressure increase
  • 194711 03/04/07 09:35 X. Zheng (TECH) Left Q3 quench
  • 194709 03/04/07 09:22 X. Zheng Delta scan finished and starting pass change
  • 194706 03/04/07 08:58 A. Camsonne Improved raster correction
  • 194697 03/04/07 08:04 A. Camsonne Fist raster corrections
  • 194693 03/04/07 07:40 A. Puckett SLI screengrab
  • 194680 03/04/07 06:09 A. Puckett scattering chamber vacuum pressure, continued
  • 194677 03/04/07 05:55 A. Puckett Scattering Chamber vacuum pressure on the high end of comfortable
  • 194670 03/04/07 05:04 A. Puckett Raster On, Spot Check
  • 194660 03/04/07 04:38 A. Puckett Controlled Access for Q3 reset
  • 194659 03/04/07 04:33 A. Camsonne BPM/raster current/position constants
  • 194658 03/04/07 03:59 A. Puckett SLI screen grab 2: Stripchart
  • 194657 03/04/07 03:58 A. Puckett SLI screen grab 1: GUI
  • 194655 03/04/07 03:52 A. Puckett target screen grab
  • 194654 03/04/07 03:48 A. Puckett Scattering Chamber Vacuum Pressure
  • 194650 03/04/07 03:11 A. Puckett Right Dipole NMR IOC reboot
  • 194639 03/04/07 00:52 A. Puckett submitted revised BTA for 3/3/07 owl shift
  • 194637 03/04/07 00:46 A. Puckett Beam Time accounting from Saturday owl shift
  • 194634 03/04/07 00:06 solvigno SLI stripchart + hall A general tools
  • 194633 03/04/07 00:02 Solvigno Shift summary
  • Saturday

  • 194624 03/03/07 22:29 Solvigno Delta scan (external sieve only in)
  • 194617 03/03/07 22:10 X. Zhan Target Screen
  • 194616 03/03/07 22:10 X. Zhan Target Screens
  • 194606 03/03/07 21:11 Solvigno Strip chart of FSI + HallA general tool
  • 194602 03/03/07 20:51 Solvigno Bull's eye scan
  • 194599 03/03/07 20:40 Solvigno FSI striptool
  • 194592 03/03/07 20:09 X. Zhan Raster Size
  • 194588 03/03/07 20:00 X. Zhan Spot++
  • 194566 03/03/07 18:37 Solvigno Good spot++ at center of target and BPM positions.
  • 194563 03/03/07 18:08 A. Camsonne Spot beam position
  • 194562 03/03/07 18:05 solvigno Spot++ at center of C target
  • 194557 03/03/07 17:43 Solvigno Aperture scan (cont'd)
  • 194543 03/03/07 16:01 X. Zheng Shift summary
  • 194540 03/03/07 15:50 X. Zheng Aperture scan run list (so far)
  • 194517 03/03/07 14:29 B. Craver target pressure spike
  • 194512 03/03/07 14:02 X. Zheng Right HRS angle (floor mark input)
  • 194511 03/03/07 13:47 X. Zheng beam vertical size
  • 194507 03/03/07 13:31 X. Zheng Good beam (almost?)
  • 194506 03/03/07 13:28 B. Craver target to BeO
  • 194505 03/03/07 12:47 B. Craver target pressure alarm
  • 194504 03/03/07 12:15 X. Zheng Right NMR coming back (slowly)
  • 194503 03/03/07 11:48 X. Zheng (TECH) Right NMR blank
  • 194501 03/03/07 06:51 A. Puckett target screen grabs, pt. 2
  • 194500 03/03/07 06:50 A. Puckett Target Screen Grabs, Pt. 1
  • 194499 03/03/07 06:40 Adaq shift status
  • Friday

  • 194497 03/02/07 22:42 solvigno More target screens
  • 194496 03/02/07 22:41 solvigno Target snapshot
  • 194495 03/02/07 22:40 S.Abrahamyan shift status
  • 194489 03/02/07 16:20 E.Chudakov Shift summary
  • 194485 03/02/07 15:49 Xiaohui Target Screen
  • 194484 03/02/07 15:09 R. Michaels (TECH) R-HRS door not working
  • 194483 03/02/07 14:47 jtewing Loss of lead flow-NMR reboot
  • 194482 03/02/07 13:50 Ole Hansen Correction
  • 194481 03/02/07 13:49 Ole Hansen L-Q1 reset
  • 194480 03/02/07 13:48 Ole Hansen Magnet problems
  • 194479 03/02/07 09:18 R. Michaels PID of R-arm: electron
  • 194478 03/02/07 07:48 jose target cool-down
  • Thursday

  • 194477 03/01/07 22:46 J.Lopez Sweep
  • 194475 03/01/07 17:14 folts tech
  • 194471 03/01/07 15:47 atrig Trigger-Download
  • 194467 03/01/07 14:23 A. Camsonne EPICS variable for OTR and SLI
  • 194446 03/01/07 01:36 A. Camsonne ROC3 power cycled
  • 0 03/01/07 01:19 no_name no_key